# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 4.1 Non-negative production strategies
The production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in Definition 4.4 are the production cost in terms of the amount required to buy the required tradables $\phi$. They do not contain any amount for the illiquid assets $\psi$, even though the illiquid assets in general reduce production cost in terms of the cost of the tradables required (Section 5.3).
@@
The terminal value $Y_{i_{max}}$ is typically zero if there are no non-zero cash flows of $\mathcal{L}$ after date $i_{max}$, specifically after terminal date $T$, and can then be disregarded. Non-zero $Y_{i_{max}}$ are needed in particular for patching together production strategies (Section 4.3).
Remark 4.7.
Production strategies are investment strategies $\phi$ with incoming and outgoing cash flows in the sense of Definition 3.1 for $A^{\prime}_{i}\geq L_{i}$, as we now show. At dates $t\in D$ with $t\neq i\in\{0,1,\ldots,T\}$, by (15), the incoming cash flows are $\tilde{Z}_{t}^{\psi+\mathcal{L}}$ and the outgoing cash flows $X_{t}^{\mathcal{L}}$. For dates $i\in\{0,1,\ldots,T\}$, to get an equation corresponding to (15), we need to link the portfolio $\phi_{i}$ to the subsequent portfolio $\phi_{\gamma(i)}$. As $A^{\prime}_{i}\geq L_{i}$, we have $A^{\prime}_{i}=L_{i}+C^{\prime}_{i}$, and we insert into this expression the definition of $A^{\prime}_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$ and $L_{i}=X_{i}^{\mathcal{L}}+\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ from Definition 4.1 and $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\phi_{\gamma(i)}\cdot S_{i}-C_{i}$ from (16) and rearrage to isolate $\phi_{\gamma(i)}\cdot S_{i}$ on the left-hand side to get:
In particular, the incoming cash flows are $\tilde{Z}_{i}^{\psi+\mathcal{L}}+C_{i}$ and the outgoing cash flows $X_{i}^{\mathcal{L}}+C^{\prime}_{i}$, i.e. with the addition of out- and subsequent inflows of capital.
A strategy according to Definition 4.4 is specified for the states with $A^{\prime}_{i}\geq L_{i}$ but may not be defined for the states with $A^{\prime}_{i}<L_{i}$. But let’s assume it is defined for this case. Inserting the definitions for $A^{\prime}_{i}$, $L_{i}$, and $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ into $A^{\prime}_{i}<L_{i}$ and rearranging terms, we get
$\phi_{\gamma(i)}\cdot S_{i}>\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}+C_{i}-X_{i}^{\mathcal{L}}$ (18)
So, when $A^{\prime}_{i}<L_{i}$, even with the capital $C_{i}$ raised, the available assets (right-hand side) are not sufficient to continue the production strategy by moving to the portfolio $\phi_{\gamma(i)}$ (left-hand side) and the strategy fails. We further discuss this in Section 4.2.
@@ 4.2 Production strategies covering failure
Production strategies $\phi$ for liabilities $\mathcal{L}$ are defined in Definition 4.4 for arbitrary fulfillment conditions, so fulfillment may only be required in sufficiently many cases ($A^{\prime}_{i}\geq L_{i}$) and not all cases as for perfect replication. For the following, we assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in $L_{i}$ has been calculated for all cases, including $A^{\prime}_{i}<L_{i}$. Proposition 8.2 in Section 8.1 in the Appendix shows that, for a positively homogeneous financiability condition, any production strategy $\phi$ for given liabilities $\mathcal{L}$ with an arbitrary fulfillment condition can be extended to a production strategy $\widetilde{\phi}$ for the almost sure fulfillment condition for adjusted liabilities $\widetilde{\mathcal{L}}$ that are identical to the non-adjusted liabilities $\mathcal{L}$ in “sufficiently many cases”. The adjusted liabilities $\widetilde{\mathcal{L}}$ can thus be seen as a “redefinition” of $\mathcal{L}$ that is “allowed” by the applicable fulfillment condition.
However, by the assumptions from Section 4.1, balance sheet insolvency $A^{\prime}_{i}<L_{i}$ implies failure, so insolvency laws take over. So, the procedure only works in reality if the “redefinition” is consistent with applicable insolvency laws. The requirement on the insolvency laws is that, in case of failure, there is in particular no bankruptcy liquidation but instead all outstanding incoming and outgoing cash flows are proportionally reduced by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities. To avoid breaking the flow, the precise definitions and analyses are provided in Section 8.1 in the Appendix.
@@ 4.3 Constructing production strategies
4.4 Definition of the value
The production cost of liabilities $\mathcal{L}$ with terminal value $Y_{T}$ for given fulfillment and financiability condition, illiquid assets $\psi$, and restrictions $\mathcal{R}^{+}$ are in general not unique, so to get a unique value, it is natural to take the essential infimum $\operatorname*{inf\leavevmode\nobreak\ ess}$ of the production cost, with the convention $\operatorname*{inf\leavevmode\nobreak\ ess}\emptyset:=+\infty$.
Definition 4.8 (Value of liabilities).
For insurance liabilities $\mathcal{L}$ with terminal value $Y_{T}$ and illiquid assets $\psi$, the value $\bar{v}_{i}^{\psi}(\mathcal{L})$ at $i\in\{0,\ldots,T-1\}$ is defined as
$\bar{v}_{i}^{\psi}(\mathcal{L}):=\operatorname*{inf\leavevmode\nobreak\ ess}\{\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\}$ (19)
where the essential infimum is taken over all production strategies $\phi$ from $i$ to $T$ with capital amounts $\mathcal{C}=(C_{j}(\phi))_{j=i}^{T}$ for given fulfillment and financiability condition and restrictions $\mathcal{R}^{+}$.
We do not study in this paper when the essential infimum above is given by a minimal production strategy.
@@ 4.5 General production strategies
Definition 4.9 below extends non-negative production strategies ϕ+italic-$\phi\in\mathcal{R}^{+}$ from Definition 4.4 under specific assumptions to what we call general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$, so also allowing for short position and with non-negative values. This extension is in particular needed for Theorem 5.3. Writing ϕitalic-$\phi\in\mathcal{R}^{\prime}$ as $\phi=\phi^{+}-\phi^{-}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$, the extension is defined by interpreting $\phi$ as the non-negative production strategy $\phi^{+}$ for the liabilities +*$\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$, with *$\mathcal{L}^{*}({\phi^{-}})$ as in Definition 3.4. This requires suitable short positions to be available for production with close out (Definition 3.3). As a simplification, we restrict to perfect production, i.e. the almost sure fulfillment condition. For more general fulfillment conditions, the liability *$\mathcal{L}^{*}({\phi^{-}})$ is potentially not fulfilled almost surely, so we would have short positions of tradables that are only fulfilled “in sufficiently many cases”. Proposition 4.10 below shows that, under the above assumptions, general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$ are characterized algebraically by the same conditions (a), (b), and (c) as non-negative production strategies in Definition 4.4.
@@
Let dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, liabilities $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, the almost sure fulfillment condition, a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$ be given. Assume that portfolios containing negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 3.3). A general production strategy $\phi$ for producing the liabilities $\mathcal{L}$ with terminal value $Y_{i_{max}}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital and illiquid assets is a strategy $(\phi)_{i=i_{min}}^{i_{max}}\in\mathcal{R}^{\prime}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that ϕ++superscriptitalic-$\phi^{+}\in\mathcal{R}^{+}$ is a non-negative production strategy with capital amounts $\mathcal{C}$ and illiquid assets $\psi$ for the sum +*$\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$ (Definition 3.4) with terminal value $Y_{i_{max}}$. The value $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for $i\in\{i_{min},\ldots,i_{max}-1\}$ is
@@
Let the almost sure fulfillment condition apply and portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 3.3). A strategy ϕitalic-$\phi\in\mathcal{R}^{\prime}$ is a general production strategy for the liabilities $\mathcal{L}$ if and only if the conditions (a), (b) and (c) from Definition 4.4 hold (formally) for $\phi$ and $\mathcal{L}$.
@@ 5 Properties of the valuation
5.1 Perfect production as extension of the market price
For this section, we assume perfect production (Definition 4.2), portfolios containing negative units in $\mathcal{R}^{\prime}$ to be available for production with close out (Definition 3.3), and illiquid assets $\psi=0$. The following Theorem 5.3 then provides necessary and sufficient conditions for the value $\bar{v}_{i}=\bar{v}_{i}^{\psi=0}$ from Definition 4.8 with general production strategies (Definition 4.9) to be an extension of the valuation of investment strategies $\phi$ of tradables by the market price $v_{i}(\phi)$ as in (4). We restrict to non-negative ϕ+italic-$\phi\in\mathcal{R}^{+}$ for simplicity. We show this by showing that a suitable liability corresponding to $\phi$ can broadly be produced with production cost of $v_{i}(\phi)$ and with no lower production cost. We do this for the liability $\mathcal{L}(\phi)$ as in Definition 3.3 with terminal value $Y_{T}=0$. As production strategies allow for capital investments, which can be seen as financial instruments in addition to tradables, it is plausible that production cost can only be an extension of the market price if the financiability condition is consistent with the tradables as in Definition 4.3.
The proof of Theorem 5.3 uses the following two lemmas. The first lemma examines the natural production strategy for $\mathcal{L}(\phi)$ given by $\phi$ itself with an adjustment for the stopping time $\tau$.
@@ 5.1 Perfect production as extension of the market price
Let the almost sure fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 3.3), and illiquid assets $\psi=0$. For ϕ+italic-$\phi\in\mathcal{R}^{+}$ on $D\equiv D[0,T]$ with zero incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}$, let $\mathcal{L}(\phi)$ be the liability from Definition 3.3. Define the strategy ϕ+superscriptitalic-$\phi^{\prime}\in\mathcal{R}^{+}$ for $t\in D$ by $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ and $\phi_{\gamma(T)}^{\prime}=0$. Then, for any $t\in D$, the outgoing cash flows $X_{t}^{\prime}$ of ϕsuperscriptitalic-$\phi^{\prime}$ are $X_{t}^{\prime}=X_{t}^{\mathcal{L}(\phi)}$, i.e. $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$. Further, the strategy ϕsuperscriptitalic-$\phi^{\prime}$ with capital $C_{i}=0$ is a production strategy for the liability $\mathcal{L}(\phi)$ with terminal value $Y_{T}=0$ and with production cost $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{t<\tau\}}$.
@@
By Definition 3.3, we have $X_{t}^{\mathcal{L}(\phi)}=1_{\{t<\tau\}}\cdot X_{t}+1_{\{t=\tau\}}\cdot(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$ for $t\in D$. We first show that the cash flows $X_{t}^{\prime}$ are given by the same expression. This holds for $t<\tau$ because then $\gamma(t)\leq\tau$, so the definition of $\phi^{\prime}$ and (2) for $\phi$ immediately imply $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{\gamma(t)}\cdot S_{t}=\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi}-X_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}$, so $X_{t}^{\prime}=X_{t}$. For $t=\tau$, we have $\gamma(t)>\tau$ and so the definition of $\phi^{\prime}$ implies $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=0=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$. For $t>\tau$, the argument is similar. Thus, $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$ for any $t\in D$. In particular, (15) from Definition 4.4 holds for $\phi^{\prime}$ for any $t\in D\setminus[0,\ldots,T]$. It further implies for $i\in[0,\ldots,T-1]$ that $A_{i+1}^{\prime}-L_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i+1}(\phi^{\prime})-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))$. For $i=T-1$, this is zero because $\phi_{\gamma(T)}^{\prime}=0$ and $\bar{v}_{T}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=Y_{T}=0$ by assumption. So the almost sure fulfillment condition and the financiability condition with $C_{i}=0$ hold. It follows that the production cost at $i=T-1$ are $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})-C_{i}=v_{i}(\phi^{\prime})$. Thus, $A_{i+1}^{\prime}-L_{i+1}=0$ at $i=T-2$. The proof proceeds recursively. ∎
@@
We show the two expressions (21) by applying consistency with the financiability condition from Definition 4.3 recursively backward in time to the strategy $\xi=\theta-\phi^{\prime}$ with $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ from Lemma 5.1. For $i\in\{0,\dots,T-1\}$, because of the almost sure fulfillment condition, $0\leq C_{i+1}^{\prime,\theta}=A_{i+1}^{\prime,\theta}-L_{i+1}^{\theta}=\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))$. Inserting the expression for $X_{i+1}^{\mathcal{L}(\phi)}$ derived from $v_{i+1}(\phi^{\prime})=\phi_{\gamma({i+1})}^{\prime}\cdot S_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}$ by Lemma 5.1, we can write this as

### v2

@@ 3.1 Non-negative production strategies
In case production strategies in Definition 3.7 must have non-negative production cost, a production strategy with zero production cost can be derived from a production strategy with negative production cost by decreasing the capital investment to $C_{i}-(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))=v_{i}(\phi)$, which by assumption is non-negative and less than $C_{i}$.
@@
The terminal value $Y_{i_{max}}$ is typically zero if there are no non-zero cash flows of $\mathcal{L}$ after date $i_{max}$, specifically after terminal date $T$, and can then be disregarded. Non-zero $Y_{i_{max}}$ (and the set $M$) are in particular needed for patching together production strategies (Section 3.4).
Value of insurance contracts. The value of insurance contracts $\mathcal{L}$ is defined as the minimum or more generally the essential infimum over the applicable production strategies. We do not study in this paper when the essential infimum is given by a minimal production strategy.
Definition 3.10 (Value of insurance contracts).
Given fulfillment and financiability condition, illiquid assets $\psi$, and restrictions $\mathcal{R}^{\geq 0}$, the value $\bar{v}_{i}^{\psi}(\mathcal{L})$ at $i\in\{0,\ldots,T-1\}$ of insurance contracts $\mathcal{L}$ with terminal value $Y_{T}$ is defined as
$\bar{v}_{i}^{\psi}(\mathcal{L})=\operatorname*{inf\leavevmode\nobreak\ ess}\{\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\}$ (19)
with the essential infimum taken over all production strategies $\phi$ from $i$ to $T$ with illiquid assets $\psi$ for the given fulfillment and financiability condition and restrictions $\mathcal{R}^{\geq 0}$. By convention, $\operatorname*{inf\leavevmode\nobreak\ ess}\emptyset:=+\infty$.
3.2 Negative production cost and future cash inflows
Negative production cost are related to the question whether future cash inflows, such as premiums or inflows from illiquid assets, can be used to pay for insurance liability cash outflows occuring at earlier dates. This is a problem that may not always be captured by solvency capital requirements. As a very simple illustrative example, we consider two one-year periods from dates $i=0$ to $i=1$, and $i=1$ to $i=2$, and deterministic insurance contract premium cash inflows $(\tilde{Z}_{1}^{\mathcal{L}},\tilde{Z}_{2}^{\mathcal{L}})=(0,100)$ and claims cash outflows $(X_{1}^{\mathcal{L}},X_{2}^{\mathcal{L}})=(10,0)$. Although the premium inflows $\tilde{Z}_{2}^{\mathcal{L}}$ significantly exceed the claims outflows $X_{1}^{\mathcal{L}}$, they can obviously not be used directly to pay the claims at date $i=1$ because they occur at the later date $i=2$. So, without additional tradables, default occurs at $i=1$. In line with this, without further assumptions, the production cost at $i=0$ is positive (e.g. equal to 10 discounted risk-free from $i=1$ to $i=0$). At the same time, the production cost at $t=1$ (which are after the cash flow $X_{1}^{\mathcal{L}}=10$) would be negative, as the cash inflows at $t=2$ of $\tilde{Z}_{2}^{\mathcal{L}}=100$ exceed the outflows of $X_{2}^{\mathcal{L}}=0$. The assumption about taking on and closing out additional liabilities set out in Section 3.1 allows “pulling back” the negative production cost at $t=1$ to $t=0$ by allowing to borrow tradables at $t=1$ with which to pay the outflows $X_{1}^{\mathcal{L}}=10$ and to subsequently pay back the borrowed tradables with the capital $C_{1}$ raised at $t=1$, for which the capital investors are compensated with the excess of the inflows $\tilde{Z}_{2}^{\mathcal{L}}=100$ over the outflows $X_{2}^{\mathcal{L}}=0$ at $t=2$.
@@ 3.3 Production strategies covering failure
Production strategies $\phi$ for liabilities $\mathcal{L}$ are defined in Definition 3.7 for arbitrary fulfillment conditions, so fulfillment may only be required in sufficiently many cases ($A^{\prime}_{i}\geq L_{i}$) and not all cases as for full replication. For the following, we assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in $L_{i}$ has been calculated for all cases, including $A^{\prime}_{i}<L_{i}$. Proposition 7.2 in Section 7.1 in the Appendix shows that, for a positively homogeneous financiability condition, any production strategy $\phi$ for given liabilities $\mathcal{L}$ with an arbitrary fulfillment condition can be extended, by covering the cases in which there is failure $A^{\prime}_{i}<L_{i}$, to a production strategy $\widetilde{\phi}$ for the full fulfillment condition for adjusted liabilities $\widetilde{\mathcal{L}}$ that are identical to the non-adjusted liabilities $\mathcal{L}$ in “sufficiently many cases”. The adjusted liabilities $\widetilde{\mathcal{L}}$ can thus be seen as a “redefinition” of $\mathcal{L}$ that is “allowed” by the applicable fulfillment condition.
However, by the assumptions from Section 3.1, balance sheet insolvency $A^{\prime}_{i}<L_{i}$ implies failure, so insolvency laws take over. So, the procedure only works in reality if the “redefinition” is consistent with applicable insolvency laws. The requirement on the insolvency laws is that, in case of failure, there is in particular no bankruptcy liquidation but instead all outstanding cash inflows and outflows are proportionally reduced by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities. To avoid breaking the flow, the precise definitions and analyses are provided in Section 7.1 in the Appendix.
@@ 3.5 General production strategies
Definition 3.11 below extends non-negative production strategies $\phi\in\mathcal{R}^{\geq 0}$ from Definition 3.7 under specific assumptions to what we call general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$, so also allowing for short position and with non-negative values. This extension is in particular needed for Theorem 4.3. Writing ϕitalic-$\phi\in\mathcal{R}^{\prime}$ as $\phi=\phi^{+}-\phi^{-}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$, the extension is defined by interpreting $\phi$ as the non-negative production strategy $\phi^{+}$ for the liabilities $\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$, with $\mathcal{L}^{*}({\phi^{-}})$ as in Definition 2.5. This requires suitable short positions to be available for production with close out (Definition 2.4). As a simplification, we restrict to full production, i.e. the full fulfillment condition. For more general fulfillment conditions, the liability $\mathcal{L}^{*}({\phi^{-}})$ is potentially not fulfilled almost surely, so we would have short positions of tradables that are only fulfilled “in sufficiently many cases”. Proposition 3.12 below shows that, under the above assumptions, general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$ are characterized algebraically by the same conditions (a), (b), and (c) as non-negative production strategies in Definition 3.7.
@@
Let dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, liabilities $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, the full fulfillment condition, a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$ be given. Assume that portfolios containing negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 2.4). A general production strategy $\phi$ for producing the liabilities $\mathcal{L}$ with terminal value $Y_{i_{max}}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital and illiquid assets is a strategy $(\phi)_{i=i_{min}}^{i_{max}}\in\mathcal{R}^{\prime}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that $\phi^{+}\in\mathcal{R}^{\geq 0}$ is a non-negative production strategy with capital amounts $\mathcal{C}$ and illiquid assets $\psi$ for the sum $\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$ (Definition 2.5) with terminal value $Y_{i_{max}}$. The value $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for $i\in\{i_{min},\ldots,i_{max}-1\}$ is
@@
Let the full fulfillment condition apply and portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 2.4). A strategy ϕitalic-$\phi\in\mathcal{R}^{\prime}$ is a general production strategy for the liabilities $\mathcal{L}$ if and only if the conditions (a), (b) and (c) from Definition 3.7 hold (formally) for $\phi$ and $\mathcal{L}$.
@@
Conditions (b) and (c) follow from showing recursively backward in time that the expression $A_{i}^{\prime}-L_{i}$ for $i\in\{0,\ldots,i_{max}\}$ for $\phi^{+}$ and $\mathcal{L}+\mathcal{L}^{*}$, which is $A_{i}^{\prime}-L_{i}=\phi_{i}^{+}\cdot S_{i}+\tilde{Z}_{i}^{\phi^{+}+\psi+\mathcal{L}+\mathcal{L}^{*}}-X_{i}^{\mathcal{L}+\mathcal{L}^{*}}-\bar{v}_{i}^{\phi^{+},\psi,\mathcal{C}}(\mathcal{L}+\mathcal{L}^{*}$), is the same expression as for $\phi$ and $\mathcal{L}$. This follows for $i=i_{max}$ and $i<i_{max}$ from Definition 2.5, the expressions for $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in Definition 3.11, and (3). ∎
@@ 4 Properties of the valuation
4.1 Valuation in the framework as extension of market price valuation
Theorem 4.3 below provides the crucial proof that, under suitable assumptions, specifically on the financiability condition, the valuation framework for liabilities introduced in this paper is an extension of the valuation of tradable investment strategies $\phi$ by the market price $v_{i}(\phi)$ as in (3). Similar to “classical no-arbitrage valuation”, the setting is full production (Definition 3.5), illiquid assets $\psi=0$, and general production strategies (Definition 3.11), and assuming that portfolios containing negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 2.4). Unlike “classical no-arbitrage valuation”, we additionally allow for capital investments. We restrict to non-negative investment strategies $\phi\in\mathcal{R}^{\geq 0}$ for simplicity and show that a suitable liability corresponding to $\phi$ can broadly be produced with production cost $v_{i}(\phi)$ and no lower production cost. We do this for the liability $\mathcal{L}(\phi)$ defined in Definition 2.4 with terminal value $Y_{T}=0$. Capital investments can be seen as financial instruments in addition to tradables, and it is shown in Theorem 4.3 that production cost can only be an extension of the market price if the financiability condition is consistent with the tradables in the sense of Definition 3.6.
The proof of Theorem 4.3 uses the following two lemmas. The first lemma examines the natural production strategy for $\mathcal{L}(\phi)$, which is $\phi$ itself with an adjustment for the stopping time $\tau$.
@@ 4.1 Valuation in the framework as extension of market price valuation
Let the full fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 2.4), and illiquid assets $\psi=0$. For a strategy $\phi\in\mathcal{R}^{\geq 0}$ on $D\equiv D[0,T]$ with zero cash inflows $\tilde{Z}_{t}=0$ and cash outflows $X_{t}$, let $\mathcal{L}(\phi)$ be the liability from Definition 2.4. Define the strategy $\phi^{\prime}\in\mathcal{R}^{\geq 0}$ for $t\in D$ by $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ and $\phi_{\gamma(T)}^{\prime}=0$. Then, for any $t\in D$, the cash outflows $X_{t}^{\prime}$ of $\phi^{\prime}$ are $X_{t}^{\prime}=X_{t}^{\mathcal{L}(\phi)}$, i.e. $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$. Further, the strategy $\phi^{\prime}$ with capital $C_{i}=0$ is a production strategy for the liability $\mathcal{L}(\phi)$ with terminal value $Y_{T}=0$, which has production cost $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{t<\tau\}}$.
@@
By Definition 2.4, we have $X_{t}^{\mathcal{L}(\phi)}=1_{\{t<\tau\}}\cdot X_{t}+1_{\{t=\tau\}}\cdot(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$ for $t\in D$. We first show that the cash flows $X_{t}^{\prime}$ are given by the same expression. This holds for $t<\tau$ because then $\gamma(t)\leq\tau$, so the definition of $\phi^{\prime}$ and (2) for $\phi$ immediately imply $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{\gamma(t)}\cdot S_{t}=\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi}-X_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}$, so $X_{t}^{\prime}=X_{t}$. For $t=\tau$, we have $\gamma(t)>\tau$ and so the definition of $\phi^{\prime}$ implies $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=0=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$. For $t>\tau$, the argument is similar. Thus, $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$ for any $t\in D$. In particular, (17) from Definition 3.7 holds for $\phi^{\prime}$ for any $t\in D\setminus[0,\ldots,T]$. It further implies for $i\in[0,\ldots,T-1]$ that $A_{i+1}^{\prime}-L_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i+1}(\phi^{\prime})-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))$. For $i=T-1$, this is zero because $\phi_{\gamma(T)}^{\prime}=0$ and $\bar{v}_{T}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=Y_{T}=0$ by assumption. So the full fulfillment condition and the financiability condition with $C_{i}=0$ hold. It follows that the production cost at $i=T-1$ are $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})-C_{i}=v_{i}(\phi^{\prime})$. Thus, $A_{i+1}^{\prime}-L_{i+1}=0$ at $i=T-2$. The proof proceeds recursively. ∎
@@
We show the two expressions (21) by applying consistency with the financiability condition from Definition 3.6 recursively backward in time to the strategy $\xi=\theta-\phi^{\prime}$ with $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ from Lemma 4.1. For $i\in\{0,\dots,T-1\}$, because of the full fulfillment condition, $0\leq C_{i+1}^{\prime,\theta}=A_{i+1}^{\prime,\theta}-L_{i+1}^{\theta}=\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))$. Inserting the expression for $X_{i+1}^{\mathcal{L}(\phi)}$ derived from $v_{i+1}(\phi^{\prime})=\phi_{\gamma({i+1})}^{\prime}\cdot S_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}$ by Lemma 4.1, we can write this as